### v1

# I4VGen: Image as Stepping Stone for Text-to-Video Generation

Abstract
Text-to-video generation has lagged behind text-to-image synthesis in quality and diversity due to the complexity of spatio-temporal modeling and limited video-text datasets. This paper presents I4VGen, a training-free and plug-and-play video diffusion inference framework, which enhances text-to-video generation by leveraging robust image techniques. Specifically, following text-to-image-to-video, I4VGen decomposes the text-to-video generation into two stages: anchor image synthesis and anchor image-guided video synthesis. Correspondingly, a well-designed generation-selection pipeline is employed to achieve visually-realistic and semantically-faithful anchor image, and an innovative Noise-Invariant Video Score Distillation Sampling is incorporated to animate the image to a dynamic video, followed by a video regeneration process to refine the video. This inference strategy effectively mitigates the prevalent issue of non-zero terminal signal-to-noise ratio. Extensive evaluations show that I4VGen not only produces videos with higher visual realism and textual fidelity but also integrates seamlessly into existing image-to-video diffusion models, thereby improving overall video quality.
Figure 1: Example results synthesized by our proposed I4VGen, a training-free and plug-and-play video diffusion inference framework, which significantly enhances the generation of visually-realistic and semantically-faithful videos.
1 Introduction
Figure 2: Illustration of non-zero terminal signal-to-noise ratio. We employ t-SNE to visualize the distributions of pure Gaussian noise, real video, and noisy video at timestep $T$, where each data point represents an independently sampled noise point or video frame. The noise schedule of AnimateDiff [25] is used, and all operations are performed in the latent space of the video autoencoder. (a) The distribution of pure Gaussian noise exhibits a disordered and diffuse nature; (b) real videos are temporally-correlated and different videos can be clearly distinguished from each other; (c) noisy videos preserve a certain degree of temporal correlation and maintain separability between different videos; (d) sampled videos for visualization.
Large-scale text-to-image models [53, 1, 52, 83, 48, 15, 22, 57, 79, 13] have demonstrated the ability to generate diverse and high-quality images from extensive web-scale image-text pair datasets. Efforts to extend these models to text-to-video synthesis [6, 25, 30, 61, 69, 73, 89, 3, 21, 2] have involved leveraging video-text pairs and temporal modeling. However, video generation remains inferior to image generation in both quality and diversity due to the complex nature of spatio-temporal modeling and the limited size of video-text datasets, often an order of magnitude smaller than image-text datasets. This disparity results in videos that lack the visual realism and textual fidelity achieved in image synthesis. This paper explores enhancing video generation by integrating robust image generation techniques, focusing on two innovative insights:
Image conditioning for text-to-video generation. Recent approaches [3, 88, 21, 6, 9, 39, 34, 82] have adopted image-assisted text-to-video generation, where an initial image generation step significantly enhances video output quality. This paradigm benefits from the strong capabilities of text-to-image models by using the generated images as detailed references for video synthesis. Despite its advantages, this method incurs high computational costs. Our approach builds on these insights but innovates by employing a training-free methodology to leverage existing text-to-image capabilities, thereby enhancing text-to-video generation performance without additional computational expense.
Zero terminal-SNR noise schedule. A prevalent issue in diffusion models is the non-zero terminal signal-to-noise ratio (SNR) [27, 41], which results from suboptimal noise schedule. This mismatch between the training phase, where residual signals persist at the terminal diffusion timestep $T$, and the inference phase, which uses pure Gaussian noise, creates a gap that can degrade model performance. As illustrated in Fig. 2, noisy videos generated under such conditions exhibit temporal correlation markedly distinct from the independent and identically distributed pure Gaussian noise.
To address these challenges, we introduce a novel training-free and plug-and-play inference framework, called I4VGen, which enhances text-to-video generation without necessitating further training or learnable parameters, and can be readily integrated into any existing video diffusion model. I4VGen restructures the video diffusion inference pipeline by using image as an intermediary in text-to-video generation, circumventing the non-zero terminal SNR issue and enhancing output quality.
Specifically, we decompose the text-to-video generation into generating an anchor image from the text prompt and then synthesizing a video from this image and the text. For the former, we design a generation-selection strategy for the anchor image, which includes synthesizing candidate images and selecting the most appropriate one based on a reward mechanism, ensuring the image aligns closely with the text prompt. For the latter, we develop a novel Noise-Invariant Video Score Distillation Sampling (NI-VSDS) method to animate the image, followed by a video regeneration, i.e., diffusion-denoising, process to refine the video.
Extensive quantitative and qualitative analyses demonstrate that I4VGen substantially improves the generation of visually-realistic and semantically-faithful videos across various text-to-video diffusion models (see Fig.1). Furthermore, our method can also be seamlessly integrated into the image-to-video diffusion models, thereby improving the temporal consistency and visual quality of the generated videos (see Fig. 6).
2 Related Work
Video Generative Models. The domain of video generation has seen significant advancements through the use of Generative Adversarial Networks (GANs) [67, 58, 65, 72, 59, 64, 16, 86, 62, 5, 60, 74], Variational Autoencoders (VAEs) [47, 40, 28], and Autoregressive models (ARs) [80, 18, 75, 33, 66, 17, 81, 84]. Despite these developments, synthesizing videos from text prompts remains challenging due to the complexities of modeling spatio-temporal dynamics. Recent innovations driven by the successes of diffusion models [31, 11, 63], which have been applied effectively in image generation [55, 48, 52, 57, 22, 1, 79, 45, 23] and audio synthesis [36, 7, 50, 37, 42], and underscore the emergence of substantial headway [32, 30, 29, 61, 4, 85, 56, 76, 6, 8, 12, 19, 9, 20, 44, 73, 87, 88, 34, 71, 15, 25, 21, 3, 26, 69, 70, 43] in research endeavors devoted to video synthesis from text input.
The foundational contributions of the Video Diffusion Model (VDM) [32] represents a milestone in leveraging diffusion models for video generation by adapting the 2D U-Net architecture used in image generation to a 3D U-Net capable of temporal modeling. Successive researches, such as Make-A-Video [61] and Imagen Video [30], expand video generation capabilities significantly. To enhance efficiency, subsequent models have transitioned the diffusion process from pixel to latent space [29, 89, 69, 4, 3, 25, 73], paralleling advancements in latent diffusion for images [55].
However, the direct generation of videos from text prompts remains intrinsically challenging. Recent approaches [3, 88, 21, 6, 9, 39, 34, 82, 54] have employed text-to-image synthesis as an intermediary step, enhancing overall performance. Despite these advancements, these methods still face the challenge of high computational training costs. In this study, we explore a novel training-free methodology aimed at bridging the existing gap in the field.
In addition, [10] (contemporary researches) introduces additional operations in the attention layer, i.e., cross-frame self-attention control, to enhance the video model. However, this necessitates modifications to the model architecture, whereas our method does not.
Signal-Leak Bias. Diffusion models are designed to generate high-quality visuals from noise through a sequential denoising process, which is consistent in both image and video diffusion models. During training, Gaussian noise corrupts the visual content, challenging the model to restore it to its original form. In the inference phase, the model operates on pure Gaussian noise, transforming it into a realistic visual content step-by-step.
Unfortunately, most existing diffusion models exhibit a disparity between the corrupted image during training and the pure Gaussian noise during inference. Commencing denoising from pure Gaussian noise in the inference phase deviates from the training process, potentially introducing signal-leak bias. For image diffusion models, [27, 41, 38] point out flaws in common diffusion noise schedules and sample steps, and propose to fine-tune the diffusion model to mitigate or eliminate the signal-leak bias during training, leading to improved results. [14] attempts to exploit signal-leak bias to achieve more control over the generated images. For video diffusion models, this issue becomes more pronounced. [77, 44] invert the retrieved video or generated low-quality to construct initial noise to alleviate the problem of signal-leak, improving inference quality. However, they suffer from limited diversity and cumbersome inference. At the same time, first-round inference of FreeInit [77] still exhibits a training-inference gap.
In contrast to existing methods, our approach utilizes images as the stepping stone for text-to-video generation. This novel pathway aims to produce visually-realistic and semantically-reasonable videos while maintaining manageable computational overheads, as detailed in Sec. 5.
3 Preliminaries
Video Diffusion Model. Aligned with the framework of Image Diffusion Models, Video Diffusion Models (VDMs) predominantly utilize the paradigm of latent diffusion models (LDMs). Unlike traditional methods that operate directly in the pixel space, VDMs function within the latent space defined by a video autoencoder. Specifically, a video encoder $\mathcal{E}\left(\cdot\right)$ learns the mapping from an input video $\mathbf{v}\in\mathcal{V},{\mathbf{v}}=\{{\mathbf{f}}^{1},{\mathbf{f}}^{2},\cdots,{\mathbf{f}}^{F}\}$ to a latent code $\mathbf{z}=\mathcal{E}\left(\mathbf{v}\right)=\{{\mathbf{z}}^{1},{\mathbf{z}}^{2},\cdots,{\mathbf{z}}^{f}\}$. Subsequently, a video decoder $\mathcal{D}(\cdot)$ reconstructs the input video, aiming for $\mathcal{D}\left(\mathcal{E}\left(\mathbf{v}\right)\right)\approx\mathbf{v}$. Typically, image autoencoder is used in a frame-by-frame processing manner instead of the video one, where $F=f$.
Upon training the autoencoder, a Denoising Diffusion Probabilistic Model (DDPM) [31] is employed within the latent space to generate a denoised version of an input latent $\mathbf{z}_{t}$ at each timestep $t$. During denoising, the diffusion model can be conditioned on additional inputs, such as a text embedding $\mathbf{c}=f_{\text{CLIP}}\left(\mathbf{y}\right)$ generated by a pre-trained CLIP text encoder [51], corresponding to the input text prompt $\mathbf{y}$. The DDPM model $\epsilon_{\theta}(\cdot)$, a 3D U-Net parametrized by $\theta$, optimizes the following loss:
$\mathcal{L}=\mathbb{E}_{\mathbf{z}\sim\mathcal{E}\left(\mathbf{v}\right),\mathbf{c}=f_{\text{CLIP}}\left(\mathbf{y}\right),\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{1}\right),t}\left[\|\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},\mathbf{c},t\right)\|_{2}^{2}\right],$ (1)
During inference, a latent variable $\mathbf{z}_{T}$ is sampled from the standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{1})$ and subjected to sequential denoising procedures of the DDPM to derive a refined latent representation $\mathbf{z}_{0}$. This denoised latent $\mathbf{z}_{0}$ is then fed into the decoder to synthesize the corresponding video $\mathcal{D}\left(\mathbf{z}_{0}\right)$.
Figure 3: Illustration of I4VGen. I4VGen is a training-free and plug-and-play video diffusion inference framework, which decomposes text-to-video generation into two stages: anchor image synthesis and anchor image-guided video synthesis. It employs a generation-selection strategy for the anchor image, synthesizing candidate images and selecting the most appropriate one based on a reward mechanism to ensure close alignment with the text prompt. Subsequently, a novel Noise-Invariant Video Score Distillation Sampling is developed to animate the image to a video, followed by a video regeneration process to refine that, thereby significantly improving the video quality.
Score Distillation Sampling. Score Distillation Sampling (SDS) [49, 68] employs the priors of pre-trained text-to-image models to facilitate text-conditioned 3D generation. Specifically, given a pre-trained diffusion model $\epsilon_{\theta}(\cdot)$ and the conditioning embedding $\mathbf{c}=f_{\text{CLIP}}\left(\mathbf{y}\right)$ corresponding to the text prompt $\mathbf{y}$, SDS optimizes a set of parameters $\phi$ of a differentiable parametric image generator $\mathcal{G}(\cdot)$ (e.g., NeRF [46]) using the gradient of the SDS loss $\mathcal{L}_{\text{SDS}}$:
$\nabla_{\phi}\mathcal{L}_{\text{SDS}}=w(t)\left(\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{c},t)-\epsilon\right)\frac{\partial\mathbf{x}}{\partial\phi},$ (2)
where $\epsilon$ is sampled from $\mathcal{N}\left(\mathbf{0},\mathbf{1}\right)$, $\mathbf{x}$ is an image rendered by $\mathcal{G}$, $\mathbf{z}_{t}$ is obtained by adding Gaussian noise $\epsilon$ to $\mathbf{x}$ corresponding to the timestep $t$ of the diffusion process, $w(t)$ is a constant that depends on the noising schedule. Inspired by this method, we proposes a Noise-Invariant Video Score Distillation Sampling strategy to efficiently harness the motion prior learned by a text-to-video diffusion model.
4 I4VGen
This section introduces I4VGen, a novel training-free and plug-and-play approach designed for enhancing the capabilities of pre-trained video diffusion models. As shown in Fig. 3, we factorize text-to-video generation into two stages: (1) anchor image synthesis to generate the anchor image $\mathbf{x}$ given the text prompt $\mathbf{y}$, and (2) anchor image-guided video synthesis to generate the video $\mathbf{v}$ by leveraging the text prompt $\mathbf{y}$ and the image conditioning $\mathbf{x}$. This section provides the detailed explanations of both stages in Sec. 4.1 and 4.2, respectively.
4.1 Anchor image synthesis
The objective in this stage is to synthesize the visually-realistic anchor image $\mathbf{x}$ that corresponds accurately to the given text prompt $\mathbf{y}$. This image serves as a foundation to provide appearance information for enhancing the performance of the subsequent video generation. As illustrated in Fig. 3, a well-designed generation-selection pipeline is employed to produce visually-realistic and semantically-faithful anchor image, which consist of candidate images synthesis and reward-based anchor image selection.
Candidate images synthesis. In contrast to generating a single image, this approach generates a set of candidate images to ensure the selection of an optimal example. Utilizing a pre-trained image diffusion model $\mathcal{D}_{\text{img}}(\cdot)$, we construct the candidate image set as follows:
${\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots,{\mathbf{x}}_{N}=\mathcal{D}_{\text{img}}\left(\mathbf{y},\mathbf{z}_{1}\right),\mathcal{D}_{\text{img}}\left(\mathbf{y},\mathbf{z}_{2}\right),\cdots,\mathcal{D}_{\text{img}}\left(\mathbf{y},\mathbf{z}_{N}\right),$ (3)
where $N$ denotes the number of candidate images, and $\mathbf{z}_{i}$ represents Gaussian noise.
Reward-based anchor image selection. With the help of the image reward model [78] $\mathcal{R}(\cdot)$, a promising automatic text-to-image evaluation metric aligned with human preferences, the candidate image with the highest reward score $s$ is selected as the anchor image $\mathbf{x}$, as defined by:
$\mathbf{x}$ $=\mathbf{x}_{i},\quad\text{where}\ i=\mathop{\arg\max}\limits_{i}s_{i}=\mathop{\arg\max}\limits_{i}\mathcal{R}\left(\mathbf{x}_{i}\right).$ (4)
The proposed generation-selection design facilitates the acquisition of a high-quality anchor image, particularly beneficial for complex text prompts (see Fig. 6). Notably, our method accommodates both user-provided and retrieved images, extending its applicability to a variety of custom scenarios, as discussed in Sec. 5.5.
4.2 Anchor image-guided video synthesis
Upon obtaining the anchor image $\mathbf{x}$, we copy it $F$ times to create an initial static video $\hat{\mathbf{v}}\in\mathcal{V},\hat{\mathbf{v}}=\{{\mathbf{x}},{\mathbf{x}},\cdots,{\mathbf{x}}\}$. The objective in this stage is to transform this static video to a high-quality video reflecting the text prompt $\mathbf{y}$. As illustrated in Fig. 3, we introduce static video animation and video regeneration.
Static video animation. A straightforward approach to animate the static video involves applying a diffusion-denoising process to transition from the static to dynamic state. However, this approach encounters a training-inference gap, as the model is trained on dynamic real-world videos but tested on static videos, leading to suboptimal motion quality due to the introduction of static priors, as discussed in Sec. 5.4.
Figure 4: Qualitative comparison. Each video is generated with the same text prompt and random seed for all methods. Our approach significantly improves the quality of the generated videos while showing excellent alignment with text prompts.
To address this limitation, we propose a novel method leveraging the motion prior from the pre-trained text-to-video diffusion model to animate static videos. Drawing inspiration from Score Distillation Sampling (SDS) as introduced in [49, 68], we develop the Noise-Invariant Video Score Distillation Sampling (NI-VSDS). Unlike vanilla SDS, which optimizes a parametric image generator, our approach directly parameterizes the static video $\hat{\mathbf{v}}$ and applies targeted optimization. The NI-VSDS loss function is defined as follows:
$\nabla_{\hat{\mathbf{v}}}\mathcal{L}_{\text{NI-VSDS}}=w(t)\left(\epsilon_{\theta}(\hat{\mathbf{v}}_{t},\mathbf{c},t)-\epsilon\right),$ (5)
where $\hat{\mathbf{v}}_{t}$ represents the noisy video at timestep $t$ perturbed by Gaussian noise $\epsilon$. Furthermore, we incorporate three strategic modifications:
  • (a) Rather than resampling the Gaussian noise at each iteration as in traditional SDS, we maintain a constant noise across the optimization, enhancing convergence speed.
  • (b) Optimization is confined to the initial stages of the denoising process, where noise levels are higher, focusing on dynamic information distillation.
  • (c) We implement a coarse-to-fine optimization strategy, evolving from high to low noise levels, specifically from timestep $T$ to $\tau_{\text{NI-VSDS}}$, where $T>\tau_{\text{NI-VSDS}}>0$. This approach stabilizes the optimization trajectory and yields superior motion quality.
The algorithmic implementation of our Noise-Invariant Video Score Distillation Sampling is detailed in Algorithm 1, which outlines the procedure for transitioning a static video into a dynamic one using the defined NI-VSDS loss. Remarkably, we only perform a single update from timestep $T$ to $\tau_{\text{NI-VSDS}}$, which requires fewer than 50 iterations, a significant reduction compared to the thousands typically necessary in SDS for text-to-3D synthesis. $\alpha$ is a scalar defining the step size of the gradient update. We empirically set $\tau_{\text{NI-VSDS}}=\mathsf{Int}(T\times p_{\text{NI-VSDS}})$.
Input: A T2V diffusion model $\epsilon_{\theta}(\cdot)$, a text prompt $\mathbf{y}$, a static video $\hat{\mathbf{v}}$, timestep $\tau_{\text{NI-VSDS}}$.
Output: A dynamic video.
1 Sampling $\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{1}\right)$; $\mathbf{c}=f_{\text{CLIP}}\left(\mathbf{y}\right)$
2 for $t$ $=$ $T,\cdots,\tau_{\text{NI-VSDS}}$ do
3       $\hat{\mathbf{v}}_{t}\leftarrow\mathsf{AddNoise}\left(\hat{\mathbf{v}},\epsilon,t\right)$
4       $\nabla_{\hat{\mathbf{v}}}\mathcal{L}_{\text{NI-VSDS}}\leftarrow w(t)\left(\epsilon_{\theta}(\hat{\mathbf{v}}_{t},\mathbf{c},t)-\epsilon\right)$
5       $\hat{\mathbf{v}}\leftarrow\hat{\mathbf{v}}-\alpha\cdot\nabla_{\hat{\mathbf{v}}}\mathcal{L}_{\text{NI-VSDS}}$
6
7return $\hat{\mathbf{v}}$
Algorithm 1 Noise-invariant Video SDS
Video regeneration. Upon animating the static video, we further enhance the video quality through a diffusion-denoising process. This phase does not suffer from the aforementioned training-inference gap, thus enabling more refined generation outcomes.
Notably, we have the flexibility to introduce noise at any timestep $\tau_{\text{re}}$, calculated as $\tau_{\text{re}}=\mathsf{Int}(T\times p_{\text{re}})$, followed by the corresponding denoising process. This strategy not only preserves the fine appearance textures but also reduces the required denoising steps, thus streamlining the video synthesis process and elevating the overall quality of the resulting video.
5 Experiments
5.1 Experimental settings
Implementation details. I4VGen is a training-free and plug-and-play strategy that can be adapted to existing text-to-video diffusion models. To ascertain the efficacy and adaptability of I4VGen, it iså applied to two well-regarded diffusion-based text-to-video models: AnimateDiff [25] and LaVie [73].
For AnimateDiff, the mm-sd-v15-v2 motion module [^1] https://github.com/guoyww/AnimateDiff, alongside Stable Diffusion v1.5, is utilized to synthesize 16 consecutive frames at a resolution of 512$\times$512 pixels for evaluation. For LaVie, the base-version [^2] https://github.com/Vchitect/LaVie is employed to generate 16 consecutive frames at 320$\times$512 pixels for evaluation. All other inference details adhere to the original settings described in [25] and [73], respectively. Notably, both AnimateDiff and LaVie possess inherent text-to-image capabilities when excluding the motion module. To avoid introducing additional GPU storage requirements, we leverage their corresponding image versions for text-to-image generation in I4VGen. For AnimteDiff, we empirically set $N$ = 16, $p_{\text{NI-VSDS}}$ = 0.4, $\alpha$ = 1, and $p_{\text{re}}$ = 1. For LaVie, we empirically set $N$ = 16, $p_{\text{NI-VSDS}}$ = 0.4, $\alpha$ = 1, and $p_{\text{re}}$ = 0.8. All experiments are conducted on a single NVIDIA V100 GPU (32 GB).
Benchmark. I4VGen is assessed using VBench [35], a comprehensive benchmark that evaluates video generation models across 16 disentangled dimensions, which is more authoritative than FVD. These dimensions provide a detailed analysis of generation quality from two overarching perspectives: Video Quality [^3] Video Quality includes 7 evaluation dimensions: Subject Consistency, Background Consistency, Temporal Flickering, Motion Smoothness, Dynamic Degree, Aesthetic Quality, and Imaging Quality. The first 5 evaluate temporal quality, and the last 2 evaluate frame-wise quality., focusing on the perceptual quality of the generated videos, and Video-Condition Consistency [^4] Video-Condition Consistency includes 9 evaluation dimensions: Object Class, Multiple Objects, Human Action, Color, Spatial Relationship, Scene, Appearance Style, Temporal Style, Overall Consistency. The first 6 evaluate semantics, the 7 and 8-th evaluate style, and the 9-th evaluates overall consistency., assessing how well the generated videos align with the provided conditions.
Table 1: VBench evaluation results per dimension. This table compares the performance of I4VGen with other counterparts across each of the 16 VBench dimensions.
Methods Subj. Cons. Back. Cons. Tem. Flick. Moti. Smo. Dyna. Degr. Aest. Qual. Imag. Qual. Obje. Class
AnimateDiff [25] 87.11 95.22 95.99 93.12 74.89 56.07 64.29 83.69
AnimateDiff + FreeInit [77] 90.45 96.57 96.89 95.66 70.17 59.25 63.51 87.55
AnimateDiff + I4VGen 95.17 97.73 98.51 96.45 57.72 64.68 66.18 92.59
LaVie [73] 91.65 96.30 98.03 95.73 71.94 59.64 65.13 91.25
LaVie + FreeInit [77] 92.32 96.35 98.06 95.83 71.11 59.41 63.89 89.13
LaVie + I4VGen 94.12 96.90 98.55 96.37 70.55 60.88 66.55 92.26
Methods Mult. Obje. Hum. Acti. Color Spat. Rela. Scene Appe. Style Tem. Style Over. Cons.
AnimateDiff [25] 22.61 90.40 81.73 31.55 45.61 24.40 24.49 25.71
AnimateDiff + FreeInit [77] 26.92 93.00 86.39 30.71 44.61 23.98 25.03 25.61
AnimateDiff + I4VGen 57.22 95.80 91.98 45.20 54.67 25.07 26.11 28.01
LaVie [73] 24.02 94.80 83.64 26.27 52.89 23.67 24.94 27.25
LaVie + FreeInit [77] 22.59 94.20 84.34 27.46 52.70 23.61 24.85 26.89
LaVie + I4VGen 32.77 96.20 88.59 33.81 55.64 24.35 25.62 27.68
5.2 Qualitative comparison
Fig. 4 presents a comparative analysis of our results against state-of-the-art counterparts using identical text prompts and random seeds. I4VGen excels in enhancing both the temporal consistency and the frame-wise quality, alongside superior alignment with the text prompts. For instance, in the case of playing guitar, AnimateDiff suffers from poor video quality, and FreeInit encounters an incomplete guitar in the middle of the video. In contrast, our method effectively addresses these issues, maintaining stable temporal consistency. Furthermore, while baseline methods struggle with accurate synthesis of all text-described components, e.g., NYC Times Square and suit, I4VGen generates videos that are visually realistic and closely aligned with the text prompts by using intermediate images with a generation-selection mechanism.
Table 2: Ablation study. Orange highlights generation-selection, while yellow highlights NI-VSDS.
Methods Subj. Cons. Back. Cons. Tem. Flick. Moti. Smo. Dyna. Degr. Aest. Qual. Imag. Qual. Obje. Class
w/o generation-selection 94.89 97.80 98.28 96.99 55.91 62.23 64.18 90.95
w/o NI-VSDS 96.47 98.82 98.99 97.56 28.24 65.17 65.52 92.66
AnimateDiff + I4VGen 95.17 97.73 98.51 96.45 57.72 64.68 66.18 92.59
Methods Mult. Obje. Hum. Acti. Color Spat. Rela. Scene Appe. Style Tem. Style Over. Cons.
w/o generation-selection 40.68 94.40 90.55 37.79 53.72 24.76 26.03 26.62
w/o NI-VSDS 62.84 94.80 91.95 47.57 55.80 24.88 25.72 27.91
AnimateDiff + I4VGen 57.22 95.80 91.98 45.20 54.67 25.07 26.11 28.01
5.3 Quantitative comparison
Objective evaluation. Following the protocols established by VBench [35], we evaluate I4VGen in terms of both video quality and video-text consistency. As detailed in Table 1, I4VGen outperforms all other approaches in temporal quality (higher background and subject consistency, less flickering, and better smoothness), frame-wise quality (higher aesthetic and imaging quality), and video-text consistency (greater semantics, style, and overall consistency). Although counterparts occasionally produce videos with more dynamic motion, they are often linked to inappropriate or excessive movements, as evidenced by the unnatural motion of a rabbit portrait shown in Fig. 4. I4VGen strikes a more effective balance between motion intensity and overall video quality, which is further verified in the user study.
Table 3: User study. I4VGen performs over other counterparts.
Methods User study
AnimateDiff [25] 6.67%
AnimateDiff + FreeInit [77] 22.67%
AnimateDiff + I4VGen 70.67%
LaVie [73] 24.00%
LaVie + FreeInit [77] 17.33%
LaVie + I4VGen 58.67%
User study. We conduct a subjective user study involving 10 volunteers with expertise in image and video processing. Participants are asked to select the most visually-appealing and semantically-faithful video from those generated by our proposed method and the state-of-the-art approaches, focusing on video quality, consistency with the text, and motion quality. Specifically, each participant has 15 questions. As shown in Table 3, I4VGen performs favorably against the other methods.
Table 4: Inference time.
Methods Time
AnimateDiff [25] 21.73s
AnimateDiff + FreeInit [77] 113.67s
AnimateDiff + I4VGen 53.78s
[…]
We also provide the inference time for a single video in Table 4, evaluated on a single NVIDIA V100 GPU (32 GB), where 50 videos are randomly generated to obtain an average inference time. Our method performs better than FreeInit.
Figure 5: Intermediate results visualization.
Figure 6: Adaptation on SparseCtrl.
5.4 Ablation study
On generation-selection strategy. We adopt a generation-selection strategy to create visually-realistic and semantically-faithful anchor images, which serve as a foundation for providing appearance information to enhance subsequent video generation performance. As shown in Table 2, highlighted in orange, compared to randomly synthesizing a single anchor image, the generation-selection strategy significantly improves the quality of the generated videos in terms of frame-wise quality and consistency with the text. Fig. 6 provides a visualization of the candidate images, where the reward-based selection mechanism eliminates unsatisfactory images, leading to better results.
On NI-VSDS. Directly applying the video regeneration process to static videos introduces static priors, resulting in suboptimal motion quality. As shown in Table 2, highlighted in yellow, while direct diffusion-denoising improves the temporal consistency of the generated videos, it severely sacrifices the motion dynamics, adversely affecting the motion style. In contrast, our method achieves an effective balance between motion intensity and overall video quality.
On video regeneration. Fig. 6 visualizes the intermediate results, demonstrating that the video regeneration process is essential for refining appearance details and optimizing motion patterns.
5.5 More applications
Figure 7: Adaptation on real image.
Adaptation on image-to-video diffusion models. I4VGen can be seamlessly integrated into existing image-to-video diffusion models by replacing the anchor images with the provided images, thereby enhancing the overall video quality. As shown in Fig. 6, integrating I4VGen into SparseCtrl [24] significantly improves the quality of the generated videos in terms of temporal consistency and appearance fidelity.
Adaptation on real image. Our method adapts to user-provided images, as shown in Fig. 7, where we use real images as anchor images, resulting in high-fidelity videos that are semantically consistent with the real images. Notably, our approach differs from vanilla image-to-video generation, as the synthesized videos are not completely aligned with the provided images. NI-VSDS is designed to animate static videos and is implemented as a spatio-temporal co-optimization.
6 Conclusion
The paper introduces I4VGen, a training-free and plug-and-play video diffusion inference framework enhancing text-to-video generation through advanced image techniques. It decomposes the video generation process into anchor image synthesis and anchor image-guided video synthesis, utilizing a generation-selection pipeline for the realistic anchor image and noise-invariant video score distillation sampling for animation. I4VGen effectively alleviates non-zero terminal signal-to-noise ratio issues and demonstrates improved visual realism and textual fidelity when integrated with existing video diffusion models.
Limitation and discussion. I4VGen improves the video diffusion model but requires more inference cost. As discussed in Sec. 5.3, the inference time of I4VGen is over double the baseline. Enhancing inference efficiency remains a future goal, with distillation techniques as a potential approach. Additionally, although our method and FreeInit are orthogonal, integrating both by replacing video regeneration with FreeInit fails to produce notable benefits.